# Pb-intercalated epitaxial graphene on SiC: Full insight into band structure and orbital character of interlayer Pb, and charge transfer into graphene

### v1

@@ I INTRODUCTION
Pb has raised particular interest as an intercalant for epitaxial graphene on SiC [18, 19, 20, 21, 22, 16, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], with the motivation of having graphene in proximity to 2D Pb, which might induce superconductivity and spin-orbit coupling effects in graphene [34, 35, 36, 37]. Pb-intercalated quasi-freestanding monolayer graphene (Pb-QFMLG) on SiC exhibits near charge-neutrality, the underlying reason for which is under active discussion [16, 26, 28]. A systematic investigation is required to disentangle the interplay of different charge transfer contributions from both the intercalated Pb and the SiC substrate to the graphene. To date, most studies on Pb-intercalated epitaxial graphene have concentrated on the atomic arrangement of the intercalated Pb layer [18, 20, 21, 22, 16, 23, 24, 25, 26, 27, 28, 31, 32].
@@ II.2 DFT calculations
DFT calculations for a (11) monolayer Pb on SiC were performed using the Perdew–Burke–Ernzerhof generalized gradient approximation exchange-correlation functional, as implemented in the Quantum ESPRESSO package [41, 42]. Spin-orbit coupling was included through the use of fully relativistic projector augmented wave pseudopotentials. A six bilayer SiC slab model was used, with one side having Pb atoms attached in the top-site position to the surface Si atoms and the other side having hydrogen atoms attached in the top-site position to the surface C atoms, with of vacuum separation between the slabs. The plane-wave cutoff energy and the -point mesh sampling density were 70 Ry and 16161, respectively. The in-plane lattice parameter for Pb was fixed to , corresponding to the experimentally measured lattice parameter for bulk 6H-SiC, while all atoms in the slab were relaxed along the -axis to a force convergence threshold of Ry/bohr. The resulting distance between the Pb atoms and the surface Si atoms was , which is in close agreement with X-ray standing waves (XSW) measurements for a Pb-QFMLG [26].
@@ III.2 Electronic band structure of interlayer Pb and its atomic arrangement relative to SiC
The Fermi surface of Pb-QFMLG integrated over a width of meV is shown in Fig. 2(d) covering the first and repeated BZs of Pb 333Here, the ARPES data obtained in half of the Fermi surface has been mirrored with respect to the direction of Pb ( direction of graphene). The Pb bands appear as nearly circular contours and can be fitted with a free electron approximation considering a Fermi momentum || Å-1 and an effective electron mass () of 5.3 times the rest electron mass (), as shown by the black dotted circles. These circular contours repeat with the SiC lattice periodicity, hence demonstrating a (11) order of the electronic structure of interlayer Pb relative to SiC [26]. These contours form an electron pocket around the point of Pb. Note that the part of this electron pocket falling in the first BZ corresponds to a repeated BZ contour and the counterpart that falls outside the first BZ boundary corresponds to the first BZ contour. According to Luttinger’s theorem, the free electron approximation gives an electron density of about cm-2 for interlayer Pb [47, 48]. Furthermore, we present the - dispersion along the direction of Pb in its first and repeated BZs, overlaid with the free electron parabolas, see Fig. 2(e). The quasi-free electron-like dispersion of the Pb bands indicates that the Pb-QFMLG heterostack exhibits Van der Waals interactions. Note that the measurements in the repeated BZ of Pb are taken at a photon energy of eV instead of eV. This was necessary in order to maximize the photoemission cross section for Pb in its repeated BZ, which significantly varies compared to the first BZ due to a different measurement geometry involving a tilt of about with respect to the normal emission orientation of the sample. The Fermi surface in the repeated BZ is shown separately in supplementary Fig. S1, revealing -band replicas around the point of graphene, corresponding to a (1010) periodicity relative to the graphene unit cell, consistent with the LEED and STM results [Fig. 1]. Furthermore, Fig. S2 in the supplementary information shows the fitted Pb band dispersion along its and directions, demonstrating a different band curvature along the respective high symmetry directions, evidently deviating from a free electron-like parabola. This indicates an anisotropic effective electron mass corresponding to the Pb bands, which is expected to have implications on charge transport measurements in the interlayer Pb.
@@ III.4 Comparison of interlayer Pb bands with DFT calculations
Fig. 4(c) shows the same ARPES dataset overlaid with the DFT calculations [solely shown in Fig. S3 (b), supplementary information], which closely capture the splitting of the electron-like interlayer band as well as the anisotropic band curvature along the and directions of Pb. DFT shows a band gap of meV around eV and || along , which closely agrees with the intensity suppression observed experimentally, marked by a yellow dotted circle in Fig. 4(b) [also discussed in Fig. 2]. Moreover, DFT shows two band branches crossing the Fermi energy along , which is also qualitatively in line with our experimental observations. Along , DFT predicts a meV band gap around eV and || , which is not obvious in the experimental data. However, there is a clear intensity suppression of the band dispersion along at eV, as also mentioned in Sec. III.2. Around the point, DFT matches the features ’3’ and ’4’ crossing the Fermi energy. Along , a relatively flat upward dispersion (’H’) at eV is observed experimentally, with DFT showing a similar feature but as split bands. Besides that, a similarly dispersing band feature is visible at eV, highlighted with the yellow dotted curve ’6’ in Fig. 4(b), which is not predicted by theory. These deviations can likely be due to the absence of graphene in the DFT calculations, which in reality is expected to have some interaction with the Pb layer underneath. DFT also underestimates the binding energy of the valence band maximum of SiC. ARPES measurements ( eV) identify it at a binding energy of eV (see Fig. S4 in the supplementary information), which is meV higher in binding energy than what DFT predicts. This discrepancy arises from the sensitivity of the DFT calculations to SiC’s H-terminated C-face, which does not fully represent the reality of a bulk substrate. Despite these limitations, the DFT calculations show a decent qualitative agreement with the interlayer Pb band structure observed experimentally.
@@ III.5 Light polarization dependence of interlayer band intensities: Orbital character
Based on the geometry of our ARPES experiments, LV polarization primarily probes the in-plane atomic orbitals of a sample, while LH polarization probes a mix of both in-plane and out-of-plane atomic orbitals. This is because the light beam does not probe the sample surface at grazing incidence but at an angle of with respect to the sample normal as shown in the central inset of Fig. 5. As a result, Fig. 5(c) suggests that the electron-like band around is predominantly due to the out-of-plane atomic orbitals. Along the direction of Pb, the out-of-plane orbital character remains dominant, though it gradually weakens towards the point, suggesting contributions from in-plane atomic orbitals. Higher statistics ARPES measurements along the direction of Pb, using both polarizations, are shown in Fig. S5 in the supplementary information. Along the direction, the color gradually shifts from blue to grey, suggesting a gradual mixing of in-plane and out-of-plane atomic orbitals. This transition continues, with the orbital character predominantly switching to in-plane around the point. Similarly, along the direction, the electron-like band exhibits a strong out-of-plane character around , which gradually decreases, transitioning into a mix of in-plane and out-of-plane orbital characters, eventually becoming predominantly in-plane around the point. Figs. 5(d) and (e) show constant energy surfaces (CESs) at eV below the Fermi level, integrated over a width of meV, obtained with LH and LV light polarizations, respectively. Fig. 5(f) illustrates the corresponding linear dichroism, clearly demonstrating that the orbital character is predominantly out-of-plane across the entire BZ, except near the point of Pb.
According to our DFT calculations, a (11) Pb monolayer on SiC exhibits a dominant -orbital character. DFT calculations (without spin-orbit coupling) showing the corresponding in-plane ( and ) and out-of-plane () atomic orbital weights, are provided in the supplementary information [Figs. S3 (c) and (d)]. The calculations reveal a orbital character for most of the Pb bands, while band features near the point of Pb display contributions from and orbitals, which closely aligns with our experimental findings. The dominance of orbital character in Pb bands is noteworthy, as it suggests potential strong interactions between the intercalated Pb atoms and the orbitals of graphene, possibly contributing to proximity effects in graphene. Notably, DFT also predicts a band just below the Fermi energy near the point of Pb [see supplementary Fig. S3 (d)], where our experimental data shows only a minor contribution from LH polarization. This discrepancy may arise from the interplay between the interlayer Pb and the overlying graphene.
@@ III.6 Charge transfer into the quasi-freestanding graphene layer
Figs. 6(a) and (b) show - cuts for Pb-QFMLG (same sample under the same measurement conditions except sample temperature) obtained at the point of graphene, perpendicular to its direction, measured with a photon energy of eV at sample temperatures of approximately K and K, respectively. From the momentum distribution curve (MDC) fitting of the Dirac cone dispersion, we find the Dirac point energy () to be meV at K, and meV at K, above the Fermi energy, corresponding to hole densities of cm-2 and cm-2, respectively. This corresponds to an increase in the -doping of Pb-QFMLG by cm-2 ( times higher), when decreasing the sample temperature from K to K. This increase in -doping can be attributed to the freezing out of -type dopants of SiC, suggesting that the near charge-neutrality of Pb-QFMLG is a combined effect of charge transfers from the intercalant Pb and the SiC substrate. Therefore, this rules out the possibility of a complete screening of charge transfer from the SiC substrate by the interlayer Pb. For reference, a room temperature ARPES measurement of the same Pb-QFMLG sample is provided in the supplementary material, where is found to be meV above the Fermi level, corresponding to a hole density of cm-2, as shown in Fig. S6. By using the same sample for these temperature dependent measurements, we can rule out any uncertainities in the carrier density of graphene associated with the initial quality of ZLG, the degree of intercalation or the formation of distinct Pb phases at the graphene/SiC interface.
@@ Acknowledgements.
- Schölzel et al. [2024] F. Schölzel, P. Richter, A. D. P. Unigarro, S. Wolff, H. Schwarz, A. Schütze, N. Rösch, S. Gemming, T. Seyller, and P. Schädlich, Large-area lead monolayers under cover: Intercalation, doping, and phase transformation, Small Structures , 2400338 (2024).
@@
- Wang et al. [2024] L.-L. Wang, S. Chen, M. Kolmer, Y. Han, and M. C. Tringides, Strain-modulated intercalated phases of Pb monolayer with dual periodicity in SiC(0001)-graphene interface, Applied Surface Science , 161572 (2024).

### v2

@@ I INTRODUCTION
Pb has raised particular interest as an intercalant for epitaxial graphene on SiC [18, 19, 20, 21, 22, 16, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], with the motivation of having graphene in proximity to 2D Pb, which might induce superconductivity and spin-orbit coupling (SOC) effects in graphene [34, 35, 36, 37]. Pb-intercalated quasi-freestanding monolayer graphene (Pb-QFMLG) on SiC exhibits near charge-neutrality, the underlying reason for which is under active discussion [16, 26, 28]. A systematic investigation is required to disentangle the interplay of different charge transfer contributions from both the intercalated Pb and the SiC substrate to the graphene. To date, most studies on Pb-intercalated epitaxial graphene have concentrated on the atomic arrangement of the intercalated Pb layer [18, 20, 21, 22, 16, 23, 24, 25, 26, 27, 28, 31, 32].
@@ II.2 DFT calculations
DFT calculations for a (11) monolayer Pb on SiC were performed using the Perdew–Burke–Ernzerhof generalized gradient approximation exchange-correlation functional, as implemented in the Quantum ESPRESSO package [41, 42]. SOC was included through the use of fully relativistic projector augmented wave pseudopotentials. A six bilayer SiC slab model was used, with one side having Pb atoms attached in the top-site position to the surface Si atoms and the other side having hydrogen atoms attached in the top-site position to the surface C atoms, with of vacuum separation between the slabs. The plane-wave cutoff energy and the -point mesh sampling density were 70 Ry and 16161, respectively. The in-plane lattice parameter for Pb was fixed to , corresponding to the experimentally measured lattice parameter for bulk 6H-SiC, while all atoms in the slab were relaxed along the -axis to a force convergence threshold of Ry/bohr. The resulting distance between the Pb atoms and the surface Si atoms was , which is in close agreement with X-ray standing waves (XSW) measurements for a Pb-QFMLG [26].
@@ III.2 Electronic band structure of interlayer Pb and its atomic arrangement relative to SiC
The Fermi surface of Pb-QFMLG integrated over a width of meV is shown in Fig. 2(d) covering the first and repeated BZs of Pb 333Here, the ARPES data obtained in half of the Fermi surface has been mirrored with respect to the direction of Pb ( direction of graphene). The Pb bands appear as nearly circular contours and can be fitted with a free electron approximation considering a Fermi momentum || Å-1 and an effective electron mass () of 5.3 times the rest electron mass (), as shown by the black dotted circles. These circular contours repeat with the SiC lattice periodicity, hence demonstrating a (11) order of the electronic structure of interlayer Pb relative to SiC [26]. These contours form an electron pocket around the point of Pb. Note that the part of this electron pocket falling in the first BZ corresponds to a repeated BZ contour and the counterpart that falls outside the first BZ boundary corresponds to the first BZ contour. According to Luttinger’s theorem, the free electron approximation gives an electron density of about cm-2 for interlayer Pb [47, 48]. Furthermore, we present the - dispersion along the direction of Pb in its first and repeated BZs, overlaid with the free electron parabolas, see Fig. 2(e). The quasi-free electron-like dispersion of the Pb bands indicates that the Pb-QFMLG heterostack exhibits Van der Waals interactions. Note that the measurements in the repeated BZ of Pb are taken at a photon energy of eV instead of eV. This was necessary in order to maximize the photoemission cross section for Pb in its repeated BZ, which significantly varies compared to the first BZ due to a different measurement geometry involving a tilt of about with respect to the normal emission orientation of the sample. The Fermi surface in the repeated BZ is shown separately in supplementary Fig. S1 [49], revealing -band replicas around the point of graphene, corresponding to a (1010) periodicity relative to the graphene unit cell, consistent with the LEED and STM results [Fig. 1]. Furthermore, Fig. S2 in the supplementary information [49] shows the fitted Pb band dispersion along its and directions, demonstrating a different band curvature along the respective high symmetry directions, evidently deviating from a free electron-like parabola. This indicates an anisotropic effective electron mass corresponding to the Pb bands, which is expected to have implications on charge transport measurements in the interlayer Pb.
@@ III.4 Comparison of interlayer Pb bands with DFT calculations
Fig. 4(c) shows the same ARPES dataset overlaid with the DFT calculations, solely shown in Fig. S3 (b) [49], which closely capture the splitting of the electron-like interlayer band as well as the anisotropic band curvature along the and directions of Pb. DFT shows a band gap of meV around eV and || along , which qualitatively aligns with the intensity suppression observed experimentally, marked by a yellow dotted circle in Fig. 4(b) [also discussed in Fig. 2]. Moreover, DFT shows two band branches crossing the Fermi energy along , which is also qualitatively in line with our experimental observations. Along , DFT predicts a meV band gap around eV and || , which is not obvious in the experimental data. However, there is a clear intensity suppression of the band dispersion along at eV, as also mentioned in Sec. III.2. Around the point, DFT matches the features ’3’ and ’4’ crossing the Fermi energy. Along , a relatively flat upward dispersion (’H’) at eV is observed experimentally, while DFT predicts a similar feature as split bands. The absence of splitting may simply be due to the broadness of feature ‘H’, making it difficult to resolve. Additionally, a similarly dispersing band feature appears at eV, highlighted by the yellow dotted curve ’6’ in Fig. 4(b), which is not captured by theory. This discrepancy likely stems from the structural model, which omits the graphene layer and the (1010) Pb superstructure relative to the graphene unit cell. Moreover, the raw spectral intensity along in Fig. 2(c) appears as a single broad feature ‘H’. Since raw data generally holds richer information [51], the apparent separation into two features in the 2D curvature might be misleading, suggesting that feature ‘6’ could actually originate from ‘H’ itself. However, the broadness of band ’H’ in the raw data also contrasts with the theory, potentially due to structural approximations in the latter.
Furthermore, DFT underestimates the binding energy of the SiC valence band maximum. ARPES measurements ( eV) identify it at eV (Fig. S4 [49]), which is meV higher than what DFT predicts. This discrepancy arises from the sensitivity of the DFT calculations to the Pb-Si distance, where the ideal (11) Pb monolayer and the absence of the graphene layer do not seem to accurately capture the interlayer charge transfer between Pb and SiC. Nonetheless, our DFT calculations qualitatively capture the key features of the experimentally observed Pb band structure, which is the primary focus of this work.
@@ III.5 Light polarization dependence of interlayer band intensities: Orbital character
Based on the geometry of our ARPES experiments, LV polarization primarily probes the in-plane atomic orbitals of a sample, while LH polarization probes a mix of both in-plane and out-of-plane atomic orbitals. This is because the light beam does not probe the sample surface at grazing incidence but at an angle of with respect to the sample normal as shown in the central inset of Fig. 5. As a result, Fig. 5(c) suggests that the electron-like band around is predominantly due to the out-of-plane atomic orbitals. Along the direction of Pb, the out-of-plane orbital character remains dominant, though it gradually weakens towards the point, suggesting contributions from in-plane atomic orbitals. Higher statistics ARPES measurements along the direction of Pb, using both polarizations, are shown in Fig. S5 [49]. Along the direction, the color gradually shifts from blue to grey, suggesting a gradual mixing of in-plane and out-of-plane atomic orbitals. This transition continues, with the orbital character predominantly switching to in-plane around the point. Similarly, along the direction, the electron-like band exhibits a strong out-of-plane character around , which gradually decreases, transitioning into a mix of in-plane and out-of-plane orbital characters, eventually becoming predominantly in-plane around the point. Figs. 5(d) and (e) show constant energy surfaces (CESs) at eV below the Fermi level, integrated over a width of meV, obtained with LH and LV light polarizations, respectively. Fig. 5(f) illustrates the corresponding linear dichroism, clearly demonstrating that the orbital character is predominantly out-of-plane across the entire BZ, except near the point of Pb.
According to our DFT calculations, a (11) Pb monolayer on SiC exhibits a dominant -orbital character. DFT calculations (without SOC) showing the corresponding in-plane ( and ) and out-of-plane () atomic orbital weights, are provided in supplementary information [Figs. S3 (c) and (d)] [49]. The calculations reveal a orbital character for most of the Pb bands, while band features near the point of Pb display contributions from and orbitals, which closely aligns with our experimental findings. Furthermore, these calculations suggest that the electron-like Pb band around hybridizes with the SiC bulk bands [16], which might explain the flattening of the dispersion around , as seen in Fig. S2 [49]. Moreover, the dominance of orbital character in Pb bands is noteworthy, as it suggests potential strong interactions between the intercalated Pb atoms and the orbitals of graphene, possibly contributing to proximity effects in graphene. Notably, DFT also predicts a band just below the Fermi energy near the point of Pb [Fig. S3 (d)] [49], where our experimental data shows only a minor contribution from LH polarization. Fig. S3 (b) [49] shows that the spectral weight of this band effectively decreases below the Fermi energy when the SOC is included. Hence, it is a subtle discrepancy, that can be attributed to the reduced complexity of our structural model.
@@ III.6 Charge transfer into the quasi-freestanding graphene layer
Figs. 6(a) and (b) show - cuts for Pb-QFMLG (same sample under the same measurement conditions except sample temperature) obtained at the point of graphene, perpendicular to its direction, measured with a photon energy of eV at sample temperatures of approximately K and K, respectively. From the momentum distribution curve (MDC) fitting of the Dirac cone dispersion, we find the Dirac point energy () to be meV at K, and meV at K, above the Fermi energy, corresponding to hole densities of cm-2 and cm-2, respectively. This corresponds to an increase in the -doping of Pb-QFMLG by cm-2 ( times higher), when decreasing the sample temperature from K to K. This increase in -doping can be attributed to the freezing out of -type dopants of SiC, suggesting that the near charge-neutrality of Pb-QFMLG is a combined effect of charge transfers from the intercalant Pb and the SiC substrate. Therefore, this rules out the possibility of a complete screening of charge transfer from the SiC substrate by the interlayer Pb. For reference, a room temperature ARPES measurement of the same Pb-QFMLG sample is provided in supplementary information [49], where is found to be meV above the Fermi level, corresponding to a hole density of cm-2, as shown in Fig. S6 [49]. By using the same sample for these temperature dependent measurements, we can rule out any uncertainities in the carrier density of graphene associated with the initial quality of ZLG, the degree of intercalation or the formation of distinct Pb phases at the graphene/SiC interface.
@@ Acknowledgements.
- Schölzel et al. [2024] F. Schölzel, P. Richter, A. D. P. Unigarro, S. Wolff, H. Schwarz, A. Schütze, N. Rösch, S. Gemming, T. Seyller, and P. Schädlich, Large-area lead monolayers under cover: Intercalation, doping, and phase transformation, Small Structures 6, 2400338 (2025).
@@
- Wang et al. [2024] L.-L. Wang, S. Chen, M. Kolmer, Y. Han, and M. C. Tringides, Strain-modulated intercalated phases of Pb monolayer with dual periodicity in SiC(0001)-graphene interface, Applied Surface Science 681, 161572 (2025).
@@
- [49] See Supplemental Material at [URL will be inserted by publisher] for additional ARPES and DFT results